OpenStudy (anonymous):

the percent composition by mass of magnesium MgBr2 (gram formula mass= 184 grams/mole) is equal to... a- 24/184 x 100 b- 160/184 x 100 c- 184/24 x 100 d- 184/160 x 100 PLEASE I REALLY NEED AN ANSWER FOR THIS QUESTION!!!

OpenStudy (anonymous):

a- 24/184 x 100

OpenStudy (anonymous):

The molar weight of magnesium bromide is about 184g. The molar weight of magnesium is about 24g, and the molar weight of bromine is about 80g.\[\%_{El}\approx\bigg(\frac{g_{El}}{g_{Total}}\bigg)100\%\]\[\%_{Mg}\approx\bigg(\frac{24g_{Mg}}{184g}\bigg)100\%\]\[\%_{Br}\approx\bigg(\frac{(2)80g_{Br}}{184g}\bigg)100\%\]

OpenStudy (anonymous):

letter a is the correct answer...